OpenStudy (anonymous):

x^6-64 facotr as a diffrence of square and cube please help

OpenStudy (anonymous):

is it -4(x+6) . my firend got that answer. didnt tell me how to do it tho

Directrix (directrix):

x^6-64 = (x^3)^2 - (8)^2 = ( x^3 + 8) * (x^3 - 8) That is the factorization into 2 squares.

Directrix (directrix):

>>is it -4(x+6) NO

Directrix (directrix):

Factor x^6-64 as difference of 2 cubes. Factoring pattern is attached.

OpenStudy (anonymous):

im going to try the cube and see what i get

OpenStudy (anonymous):

x--4(x^3+4x+16)

OpenStudy (anonymous):

i think

Directrix (directrix):

x^6-64 = [ (x^2)^3 - (4)^3) ] = [ x^2 - 4 ] * [ (x^2)^2 + (x^2)*4) + (-4)^2 ]

OpenStudy (anonymous):

(x-4)(x^2+4x+4^2)

Directrix (directrix):

[ x^2 - 4 ] * [ (x^2)^2 + (x^2)*4) + (-4)^2 ] = (x^2 -4) (x^4 + 4x^2 + 16)

OpenStudy (anonymous):

okay i see what i did wrong . i dont know where i got a 3 from. thank you for the help . both of you

Directrix (directrix):

I am uncertain about the instructions. Taking it from the top: x^6-64 = (x^3)^2 - (8)^2 = ( x^3 + 8) * (x^3 - 8) = (x + 2) ( x^2 -2x + 4) (x - 2) (x^2 + 2x + 4)

Directrix (directrix):

You are welcome.